OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and (1, –3) in slope-intercept form

OpenStudy (anonymous):

y = m x + b m = delta y / delta x m = y1 -y0/ x1 - x0 m = 4 - (-3) / 2 - 1 m = 7/1 m = 7 y = 7 x + b 4 = 7 * 2 + b 4 = 14 + b b = - 10 y = 7 x - 10